OpenStudy (anonymous):

I have to use the sequence defined by: \[x_{n+1} = \frac{ 1 }{ 4-x_n }\] and \[x_1 = 3\] in a proof, but before that I need to show this sequence converges. I assume the way to do this is (given where we are at in the book) to show that this sequence is monotone, and that is where I am stuck. I've done some test values, and it appears to be decreasing, how do I prove this?

OpenStudy (anonymous):

Any suggestions?

OpenStudy (anonymous):

Ok so, I tried using induction. So I wanted to show\[x_{n+1} \le x_n\] so \[\frac{ 1 }{ 4-x_n } \le x_n\] manipulating the inequality gave me a quadratic: \[x _{n}^2-4x_n+1 \le 0\]

OpenStudy (anonymous):

Do I need to apply the quadratic formula or something or am I completely off base?

OpenStudy (anonymous):

So that gives me roots at 2+/- sqrt(3)